OpenStudy (anonymous):

What is the absolute value of 11+4i?

OpenStudy (anonymous):

same as magnitude?

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\] just like pythagoras

OpenStudy (anonymous):

yeah magnitude. same thing

OpenStudy (anonymous):

no i's not minus signs, just straight up \[\sqrt{a^2+b^2}\]